a 1.555g piece of gold is warmed to 14.7 degree celsius by 16.47j of energy. if gold has a specific heat capacity of 129j/kgC what was the initial temperature of gold


1
Expert's answer
2021-02-28T06:18:24-0500

The amount of heat required to heat gold can be calculated using the formula:

"Q = C\times m\times∆t"

C=129j/kgC

Q=16.47j 

m=1.555=0.0016kg

t2=14.7

"\Delta t=\frac{Q}{C\times m}"

"(14.7-t1)=\frac{16.47}{129\times 0.0016}"

14.7-t1=79.79

t1=-65.09